9:20 SPARE MOTOR TESTING AFTER NEW CABLING

We proceed to testing the spare motor with the new cabling completed last night by Mike W.

Functional test: the motor was connected to the driver. Forward and backward spin OK. Brakes OK. Hence, new cabling works!

We now proceed to install in the W unit the spare motor to drive the y-stage.

y-motor current requirements tests

VERTICAL DISPLACEMENT TESTS

All tests below will be done moving the y-stage vertically (up and down) using the new y-motor.

UP AND DOWN +\- 40 mm:
  • ystage_sparemotor_updown_40mm.PNG:
    ystage_sparemotor_updown_40mm.PNG

UP AND DOWN +\- 80 mm:
  • ystage_sparemotor_updown_80mm_take2.PNG:
    ystage_sparemotor_updown_80mm_take2.PNG
  • ystage_sparemotor_updown_80mm.PNG:
    ystage_sparemotor_updown_80mm.PNG

HORIZONTAL DISPLACEMENT TESTS

All tests below will be done moving the y-stage horizontally (back and forth) using the new y-motor.

BACK AND FORTH +\- 80 mm:
  • ystage_hor_updown_80mm.PNG:
ystage_hor_updown_80mm.PNG
  • ystage_hor_updown_80mm_take2.PNG:
ystage_hor_updown_80mm_take2.PNG

SUMMARY OF CURRENT REQUIREMENTS

The following table reports the mean value of the currents required to move the y-stage with the new y-motor (i.e. spare motor with new cabling). All values show that the motor is operating properly.

UP DOWN FORTH BACK Notes
y-motor +1.4A +0.3A +0.4A -0.4A spare motor with new cabling

x-motor current requirements tests

The x-motor was reinstalled on the W unit, driving the x-stage as before the intervention. Just to verify that this motor is operating properly and no damage has been done, we performed the following tests.

VERTICAL DISPLACEMENT TESTS

UP AND DOWN +\-110 mm:
  • xstage_ver_updown_110mm.PNG:
    xstage_ver_updown_110mm.PNG
  • xstage_ver_updown_110mm_take2.PNG:

    xstage_ver_updown_110mm_take2.PNG


HORIZONTAL DISPLACEMENT TESTS

BACK AND FORTH +\- 40 mm:
  • xstage_hor_updown_40mm.PNG:

    xstage_hor_updown_40mm.PNG

BACK AND FORTH +\- 110 mm:
  • xstage_hor_updown_110mm.PNG:
    xstage_hor_updown_110mm.PNG

SUMMARY OF CURRENT REQUIREMENTS

The following table reports the mean value of the currents required to move the x-stage with the x-motor. All values show that the motor is operating properly as before.
UP DOWN FORTH BACK Notes
x-motor +2.2A +0.5A +0.4A -0.4A x-motor on the x-stage as before the intervention

11:00 FULL CHECK OF MOTORS/STAGES DRIVEN FROM WFS SOFTWARE

All connections for driving the x- and y- motors from the WFS software were re-established, and a full operational test driving x- and y- stages back-and-forth, and up-and-down was done successfully. This concludes the activities regarding the motor repairment!

11:45 LIMIT SWITCH CABLING REPAIRMENT

An invalid state in which BOTH limit switches of the y-stage were signaled at the same time was detected. This anomalous behavior is due to a problem with the cable supplying power to the photo-diodes of the limit switches. Mike W removed the electronics carter and detected a loose connection of the power cable. He re-solded the pin to the cable. After this repairment we verified that the invalid state does not occur anymore. Limit switches OK !!!
  • 2013-08-24_11.16.25.jpg:
    2013-08-24_11.16.25.jpg

CUBE VIGNETTING CHECK su W#1

Controlliamo le immagini di Lucifer con il cubo in varie posizioni (sistema SX)

Prendiamo le coordinate del rettangolo esterno all'immagine sfuocata.

File Posizione cubo mm Dimensione vignettamento
ima4_a Parked -68 102 px
ima3_a Nominale -4.4 323 px
Deduciamo:

scala in px/mm = (323-102)/(68-4.4) = 3.5 px/mm

Il vignettamento da parked è 102/3.5 = 29 mm

Poiché il limite del cubo W#2 è -80 mm, abbiamo solo 12 mm di corsa residua che non sono sufficienti.

NOTA: il nostro modello di stage Micos dovrebbe avere almeno 102 mm di corsa, dei quali solo 80 sono usati per non tonfare il mercury contro la flangia. Se facessimo un'apertura in questa flangia, potremmo sfruttare altri 22 mm di corsa, che sommati ai 12 residui ci porterebbero a più di 30mm dalla posizione parked, e quindi fuori dal campo di Lucifer.

Da un sopralluogo sul W#1 vediamo che andando indietro oltre -78mm la base della giostrina va a collidere con la pennina e quindi non è possibile rimuovere completamente il vignetting di lucifer semplicemente aumentando il ange dello translation stage


2013-08-24_15.46.03.jpg:
2013-08-24_15.46.03.jpg

18:00 W#2 Functional check

Non si vede ancora luce della pennina sul W#2. Saliamo e verifichiamo che il cubo non ha nessuna protezione ed è correttamente girato verso il W e il W è in asse.

Non possiamo verificare se la pennina fa luce perche' il probe di Jesper è parcheggiato in asse (COME SEMPRE!)

Dopo varie peripezie e scambi di computer AGW, il probe viene tolto. In asse non si vede assolutamente niente.

Apriamo la fioriera: lampada OK, fibra OK, dalla pennina non esce niente. Rimettendo la fibra a posto, arriva la luce.

Deduzione: la fibra era montata male e si e' spostata durante il trasporto dell'AGW...

Richiudiamo tutto e facciamo un functional test completo.

Tutti i devices OK.


SUMMARY ------------------------------------
Bin: 1 - Framerate 990 Hz (2500 kpix/sec) - RON (e-): 7.44 (estimated)
Bin: 2 - Framerate 990 Hz ( 800 kpix/sec) - RON (e-): 5.18 (estimated)
Bin: 2 - Framerate 625 Hz ( 380 kpix/sec) - RON (e-): 3.90 (estimated)
Bin: 2 - Framerate 400 Hz ( 380 kpix/sec) - RON (e-): 4.00 (estimated)
Bin: 3 - Framerate 200 Hz ( 380 kpix/sec) - RON (e-): 4.34 (estimated)
Bin: 4 - Framerate 300 Hz ( 335 kpix/sec) - RON (e-): 4.42 (estimated)
Bin: 4 - Framerate 200 Hz ( 335 kpix/sec) - RON (e-): 4.65 (estimated)
Bin: 4 - Framerate 105 Hz ( 335 kpix/sec) - RON (e-): 5.23 (estimated)

Valori da primato! Tutto il WFS e' acceso (motori slitte spente), ma il telescopio e' per la maggior parte a riposo e forse questo ci favorisce.

-- FernandoQuiros - 24 Aug 2013

  • w2_ccd39.png:
    w2_ccd39.png

  • w2_wobble.png:
    w2_wobble.png
I Attachment Action Size Date Who Comment
2013-08-24_11.16.25.jpgjpg 2013-08-24_11.16.25.jpg manage 1 MB 24 Aug 2013 - 19:13 UnknownUser  
2013-08-24_15.46.03.jpgjpg 2013-08-24_15.46.03.jpg manage 1 MB 25 Aug 2013 - 16:23 UnknownUser  
w2_ccd39.pngpng w2_ccd39.png manage 69 K 25 Aug 2013 - 03:24 AlfioPuglisi  
w2_wobble.pngpng w2_wobble.png manage 4 K 25 Aug 2013 - 03:30 AlfioPuglisi  
xstage_hor_updown_110mm.PNGPNG xstage_hor_updown_110mm.PNG manage 70 K 24 Aug 2013 - 18:26 UnknownUser  
xstage_hor_updown_40mm.PNGPNG xstage_hor_updown_40mm.PNG manage 59 K 24 Aug 2013 - 18:26 UnknownUser  
xstage_ver_updown_110mm.PNGPNG xstage_ver_updown_110mm.PNG manage 97 K 24 Aug 2013 - 18:26 UnknownUser  
xstage_ver_updown_110mm_take2.PNGPNG xstage_ver_updown_110mm_take2.PNG manage 75 K 24 Aug 2013 - 18:26 UnknownUser  
ystage_hor_updown_80mm.PNGPNG ystage_hor_updown_80mm.PNG manage 65 K 24 Aug 2013 - 18:03 UnknownUser  
ystage_hor_updown_80mm_take2.PNGPNG ystage_hor_updown_80mm_take2.PNG manage 65 K 24 Aug 2013 - 18:03 UnknownUser  
ystage_sparemotor_updown_40mm.PNGPNG ystage_sparemotor_updown_40mm.PNG manage 84 K 24 Aug 2013 - 17:40 UnknownUser  
ystage_sparemotor_updown_80mm.PNGPNG ystage_sparemotor_updown_80mm.PNG manage 70 K 24 Aug 2013 - 17:40 UnknownUser  
ystage_sparemotor_updown_80mm_take2.PNGPNG ystage_sparemotor_updown_80mm_take2.PNG manage 62 K 24 Aug 2013 - 17:40 UnknownUser  
Topic revision: r5 - 25 Aug 2013, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback